using logarithms solve (321,000)^2/7
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(57967)

using logarithms solve (321,000)^2/7
-----------------------------
Let x = (321,000)^(2/7)
---
log(x) = (7/2)log(321000)
----------------------------
log(x) = (7/2)*5.5065050032
---
log(x) = 19.27276761
---------
x = 10^19.27276761..
---
x = 1.873991486 x 10^19
==============================
Cheers,
Stan H.

Answer by Alan3354(31519)